OpenStudy (anonymous):

Angle H and angle J are complementary. Angle H measures x + 5 and angle J measures 2x – 5. Find the measure of angle H.

OpenStudy (anonymous):

Complementary = 90 degrees so H + J = 90 degrees (x+5) +(2x - 5) = 90 degrees 3x = 90 degrees x= 30 degrees So if H = x+5 H = 30 +5 = 35 degrees